# AdapEdit: Spatio-Temporal Guided Adaptive Editing Algorithm for Text-Based Continuity-Sensitive Image Editing

### 13 Dec 2023 v1

@@ 1 Introduction
Text-conditioned image synthesis has recently received tremendous success with various magnificent and creative text-to-image cases being produced, where diffusion probability models (DPMs) (Ho, Jain, and Abbeel 2020; Nichol and Dhariwal 2021), as the backbone network that plays an important role, has been developed into a variety of text-to-image generation applications (Saharia et al. 2022; Ramesh et al. 2022; Ma et al. 2023). However, the controllability of image synthesis and the support for user-customization are still challenging issues and have received more and more researchers’ attention. To solve these problems, Stable Diffusion (Rombach et al. 2022) and ControlNet (Zhang and Agrawala 2023), as two leading studies, have proposed two efficient text-conditioned learning frameworks for training, and subsequently, a series of text-guided image editing methods (Hertz et al. 2022; Ruiz et al. 2023; Wallace, Gokul, and Naik 2023; Kawar et al. 2023; Mokady et al. 2023) have been proposed, which achieve user-customized image output by editing or modifying the original image given the human-written instructions, and have witnessed a huge improvement in terms of controllability.
@@ 2 Related Work
Diffusion-Based Generative Models. Recent years has witnessed the remarkable success of diffusion-based generative models, due to their excellent performance in the diversity and impressive generative capabilities. These previous efforts mainly focus in sampling procedure (Song et al. 2020; Liu et al. 2022), conditional guidance (Dhariwal and Nichol 2021; Nichol et al. 2021), likelihood maximization (Kingma et al. 2021; Kim et al. 2022) and generalization ability (Kawar et al. 2022; Gu et al. 2022) and have enabled state-of-the-art image synthesis.
@@ 3.2 AdapEdit Algorithm Framework
Specifically, we first propose a flexible word-level temporal (FWT) adjustment module to support different guidance scales being adaptively assigned to sequential words for temporal-guided editing, which is driven by the impacts of each word $\bm{e}^{*}$ in the editing instruction $\bm{c^{*}}$ is variational. For further explanation, take the Figure 4(a) as an example, when only editing the attention map from $\textbf{M}_{t}^{\texttt{[standing]}}$ into $\textbf{M}_{t}^{\texttt{[sitting]}}$, the attention map from “dog” still remains the salient features of “standing”, which hampers from the effective editing of the above instruction “a dog sitting…” To address this issue, we design a soft attention strategy to calculate the $\tau_{\bm{c}^{*}}$ for fine-grained word-level editing in FWT module. Further, we subsequently design a dynamic pixel-level spatial (DPS) weighting module to adaptively integrate the edited visual features into original image for spatial-guided editing. Note the spatial-guidance scales $\textbf{A}_{\texttt{[S]}}$ are determined by calculating the semantic similarities between the replaced words’ average representations and the visual features. Based on the above two modules, AdapEdit can reliably control the DPMs (e.g., SD-v1.4) to effectively perform continuity-sensitive soft editing tasks with as many details as possible preserved from the original image x. Next, we detail the separate modules.
@@ 4.2 Qualitative Evaluation
The qualitative evaluation results are shown in Figure 5. From Figure 5, we obtain the following three observations. 1) For the hard editing instruction from “cake” to “chocolate cake”, all the methods show excellent performance, which is relatively easier to be achieved for diffusion-based methods. Whereas when editing “boy” into “fat boy”, previous methods basically lost the original image’s details such as the background of the “bench” and the posture of the “legs”. Even worst, InstructPix2Pix fail to editing the “boy” into “fat boy”. Compared with them, AdapEdit achieves relatively optimal editing results and retains almost lossless original characteristics. 2) Further, when editing the posture of “cat” in the second line of Figure 5 and the number of “peach” in the third line of Figure 5, Prompt-to-Prompt, SDEdit, and InstructP2P all fail to accomplish the posture- or number-oriented editing task. Benefiting from the adaptive editing capability of the FWT and DPS modules, our approach effectively copes with all hard or soft editing scenes. Moreover, as shown in Figure 2, AdapEdit can also perform adaptive editing to control that the “two apples” have different sizes with certain reasoning capability emerging. 3) MasaCtrl achieves excellent object editing performance but fails to retain more details from the original image. It is worth noting that both of these are well addressed by our AdapEdit, which proves its effectiveness.
@@ 5 Conclusion
6 Acknowledgments
This work was supported by the National Key Research and Development Program of China (2022ZD0160603), and the Project funded by China Postdoctoral Science Foundation (No. 2023M741950).

### 24 Dec 2023 v2

@@ 1 Introduction
Text-conditioned image synthesis has recently received tremendous success with various magnificent and creative text-to-image cases being produced, where diffusion probability models (DPMs) (Ho, Jain, and Abbeel 2020; Nichol and Dhariwal 2021), as the backbone network that plays an important role, has been developed into a variety of text-to-image generation applications (Saharia et al. 2022; Ramesh et al. 2022; Ma et al. 2023a). However, the controllability of image synthesis and the support for user-customization are still challenging issues and have received more and more researchers’ attention. To solve these problems, Stable Diffusion (Rombach et al. 2022) and ControlNet (Zhang and Agrawala 2023), as two leading studies, have proposed two efficient text-conditioned learning frameworks for training, and subsequently, a series of text-guided image editing methods (Hertz et al. 2022; Ruiz et al. 2023; Wallace, Gokul, and Naik 2023; Kawar et al. 2023; Mokady et al. 2023) have been proposed, which achieve user-customized image output by editing or modifying the original image given the human-written instructions, and have witnessed a huge improvement in terms of controllability.
@@ 2 Related Work
Diffusion-Based Generative Models. Recent years has witnessed the remarkable success of diffusion-based generative models, due to their excellent performance in the diversity and impressive generative capabilities. These previous efforts mainly focus in sampling procedure (Song et al. 2020; Liu et al. 2022), conditional guidance (Dhariwal and Nichol 2021; Nichol et al. 2021), likelihood maximization (Kingma et al. 2021; Kim et al. 2022) and generalization ability (Kawar et al. 2022; Ma et al. 2023b) and have enabled state-of-the-art image synthesis.
@@ 3.2 AdapEdit Algorithm Framework
Specifically, we first propose a flexible word-level temporal (FWT) adjustment module to support different guidance scales being adaptively assigned to sequential words for temporal-guided editing, which is driven by the impacts of each word $\bm{e}^{*}$ in the editing instruction $\bm{c^{*}}$ is variational. For further explanation, take the Figure 4(a) as an example, when only editing the attention map from $\textbf{M}_{t}^{\texttt{[standing]}}$ into $\textbf{M}_{t}^{\texttt{[sitting]}}$, the attention map from “dog” still remains the salient features of “standing”, which hampers from the effective editing of the above instruction “a dog sitting…”.
To address this issue, we design a soft attention strategy to calculate the $\tau_{\bm{c}^{*}}$ for fine-grained word-level editing in FWT module. Further, we subsequently design a dynamic pixel-level spatial (DPS) weighting module to adaptively integrate the edited visual features into original image for spatial-guided editing. Note the spatial-guidance scales $\textbf{A}_{\texttt{[S]}}$ are determined by calculating the semantic similarities between the replaced words’ average representations and the visual features. Based on the above two modules, AdapEdit can reliably control the DPMs (e.g., SD-v1.4) to effectively perform continuity-sensitive soft editing tasks with as many details as possible preserved from the original image x. Next, we detail the separate modules.
@@ 4.2 Qualitative Evaluation
The qualitative evaluation results are shown in Figure 5. From Figure 5, we obtain the following three observations. 1) For the hard editing instruction from “cake” to “chocolate cake”, all the methods show excellent performance, which is relatively easier to be achieved for diffusion-based methods. 2) Further, when editing the posture of “cat” in the second line of Figure 5 and the number of “peach” in the third line of Figure 5, Prompt-to-Prompt, SDEdit, and InstructP2P all fail to accomplish the posture- or number-oriented editing task. Benefiting from the adaptive editing capability of the FWT and DPS modules, our approach effectively copes with all hard or soft editing scenes. Moreover, as shown in Figure 2, AdapEdit can also perform adaptive editing to control that the “two apples” have different sizes with certain reasoning capability emerging. 3) MasaCtrl achieves excellent object editing performance but fails to retain more details from the original image. It is worth noting that both of these are well addressed by our AdapEdit, which proves its effectiveness.
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Key Research and Development Program of China (2022ZD0160603), the Project funded by China Postdoctoral Science Foundation (No. 2023M741950), and the Nationally Funded Postdoctoral Researcher Program (GZB20230347).